A concert audience of 400 people consists of adults, students, and children. The ticket prices are $40 for adults, $20 for students, and $10 for children. The total amount of money taken in was $10,600. The numbers of children tickets sold is 200 less than the number of adults and student tickets in total. How many adults, students, and children are in attendance?
Answer by ikleyn(53213)

.
A concert audience of 400 people consists of adults, students, and children. The ticket prices are $40 for adults, $20 for students,
and $10 for children. The total amount of money taken in was $10,600. The numbers of children tickets sold is 200 less
than the number of adults and student tickets in total. How many adults, students, and children are in attendance?
~~~~~~~~~~~~~~~~~~~~~~~~

Let A = # of adults tickets, S = # of students tickets.

Then the number of children tickets is A + S - 200.

You have these equations

  A +   S +    (A+S-200) =   400,    (1)
40A + 20S + 10*(A+S-200) = 10600.    (2)


Simplify and write in the standard form

 2A +  2S =   600,                  (1')
50A + 30S = 12600.                  (2')


Or even simpler

  A + S   = 300,                    (1'')
5A + 3S   = 1260                    (2'')


From (1''), express A = 300-S and substitute it into (2''). You will get

5*(300-S) + 3S = 1260.


From this point, please complete the solution on your own.